VIRGO commissioning progress
Download
1 / 54

VIRGO commissioning progress - PowerPoint PPT Presentation


  • 119 Views
  • Uploaded on
  • Presentation posted in: General

VIRGO commissioning progress. Paolo La Penna for the VIRGO collaboration European Gravitational Observatory. Virgo shutdown had started after C7 (19 September 2005) The new IB had been installed in the IB tower (end November 2005 The fake dihedron was still mounted on the bench

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'VIRGO commissioning progress ' - parker


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Virgo commissioning progress

VIRGO commissioning progress

Paolo La Penna

for the VIRGO collaboration

European Gravitational Observatory


Status in december 2005

Virgo shutdown had started after C7 (19 September 2005)

The new IB had been installed in the IB tower (end November 2005

The fake dihedron was still mounted on the bench

The Nd:YAG beam (up to 5 W for the moment) had been sent into the bench: some light (with even some interference) was visible at the end of the MC

The optics on the bench were being realigned using the IB actuators: this operation was performed staying inside the tower, with the laminar flux on and the LC closed

The alignment of the IMC was going on

(http://www.ego-gw.it/ILIAS-GW/WP1docs/lapenna_121205.ppt)

Status in December 2005


Since then

The IMC was fully aligned in vacuum at about mid December.

IMC locked at end of 2005.

Beam at the end of the ITF at begin of January 2006.

ITF alignment at mid of January.

First ITF cavities lockings at end of January.

Full RFC alignment at end of February.

Full locking of the ITF going on now (mid of March).

Since then:

The full restart of the ITF, after IB installation,

took about three months



Operation performed on the injection system

Alignment and locking of the IMC in vacuum

Alignment of the IB optics in vacuum

Alignment of the beam on the ITF and telescopes tuning

Alignment and locking of the RFC

Setup and alignment of the EIB

BMS

IMC matching improvement

IMC automatic alignment

Power stabilization

RFC automatic alignment

Operation performed on the Injection System



Injection system
Injection system

Lower stage

Upper stage


Difference old new ib
Difference old-new IB

Old IB

New IB


Imc alignment december 2005
IMC alignment: december 2005

HeNe autocollimation on the MC mirror

Acting on the IB alignment

(checked with a ccd)


Imc alignment december 20051
IMC alignment: december 2005

laser

YAG beam direct beam centering

on the MC mirror acting on the

external steering mirrors


Imc alignment december 20052
IMC alignment: december 2005

laser

MC mirror alignment

until IMC resonances are visible


Imc locking december 2005
IMC locking: december 2005

laser

Pout 8 W

IMC locking

Feedback on laser pzt

using the reflection as a trigger



Rfc alignment and locking february 2006

The Reference Cavity alignment has taken more time than expected

A first attempt to align it using the 20 W beam failed

A second attempt after having aligned a probe beam coming from the opposite direction, and looking at the RFC transmission, led to a poor RFC alignment (but it allowed to lock and start working on the ITF)

A third attempt overlapping the probe beam and the 20 W beam was successful

From the beginning to the end almost one month was necessary

RFC alignment and locking: february 2006


Yag from inverse path fringes
YAG from inverse path: fringes expected

  • Probe beam from the inverse path

  • Probe beam superposed to a HeNe passing through the RFC in the lower part

  • Camera looking at the coils frame (hit by the RFC transmission): mode flashing visible

Secondary YAG beam

1 V @ 25 Hz pzt scanning

ccd

  • CCD and PhD looking at RFC reflection

ccd

photodiode


Contrast improvement and alignment
Contrast improvement and alignment expected

YAG probe Reflected beam

Contrast = 0.7




Near field of m13 and far field
Near Field of M13 and Far Field expected

1m-f

FF

NF

Screens


M14 movement
M14 movement expected

1m-f

FF

Move M14 and observe overlapping with NF ccd

NF


M13 movement
M13 movement expected

1m-f

FF

Correct with M13 to overlap on the FF ccd

Move M14 and observe overlapping with NF ccd

NF



The rfc transmission is not visible with the rfc aligned
The RFC transmission is not visible with the RFC aligned expected

  • Inverse path beam and direct beam are being superposed below acting on M13 and M4 mirrors

  • When superimposed both beams should resonate

ccd

PhD


Rfc locked reflected 20 w yag beam 1 march 2006
RFC locked, reflected 20 W YAG beam: 1 March 2006 expected

Good alignment

+

longitudinal mismatching

Contrast = 54 %

(it was 80% before the shutdown)


Rfc alignment conclusion

The RFC alignment was more difficult than expected: the cavity axis was not where it was expected, maybe the RFC was not aligned inside its tank, or it had moved slightly during new IB assembling

The RFC is locked in a stable way, with the beam transmitted by the IMC: more stable with the alignment with respect to the previous setup

We cannot see the RFC transmission

RFC alignment: conclusion


Further activities on rfc

RFC locking filter change: a new filter has been implemented, which allows a reduction of the frequency noise of the prestabilized beam

RFC automatic alignment: this activity is going on; it will allow to keep the beam aligned on the RFC using the two closed loop pzt placed before the RFC. This allows shifts of the beam on the upper part, in order to better align the upper part optics, without losing the RFC alignment

Further activities on RFC


Problems encountered
Problems encountered implemented, which allows a reduction of the frequency noise of the prestabilized beam

Several problems/failure have occurred:

it has been necessary to open the IB tower twice

Wrong PSD

electronics

DPS misaligned

TS damaged

Beam clipped

PM stuck

RFC_trans clipped


Further activities involving the injection system

Beam Steering System implemented, which allows a reduction of the frequency noise of the prestabilized beam

IMC matching improvement

IMC automatic alignment restarted

Input power stabilization (the photodiode measuring the IMC transmission has been changed)

RFC automatic alignment

PSD on the IB working (new homemade electronics)

Further activities involving the injection system

After the restart of the IB, some new features and other activities have been performed:


Bms dc quadrants abp actuation
BMS: DC quadrants, ABP actuation implemented, which allows a reduction of the frequency noise of the prestabilized beam

  • BMS allows to keep beam aligned with respect to DC_quadrants placed on the External Injection Bench.

  • The signal from the DC quadrants is sent to the ABP (Automatic Beam Positioning) mirrors (pzt steered mirrors).

  • It allows a stable and repetitive positioning of the beam entering into the IMC;

  • It allows performing pure translations and angle of the beam.

ABPs

NF QF

FF QF

(on laser

bench)


Imc matching

With the help of BMS, IMC matching has been improved: implemented, which allows a reduction of the frequency noise of the prestabilized beam

96% matching (it was 82%),

91% contrast (should correspond to about 1500 ppm losses in the IMC).

IMC transmission is now about 8 W (at about the same level as before the shutdown)

This power is entering into the ITF (we had 0.7 W entering into the ITF before the shutdown)

IMC matching


Imc automatic alignment

The AA of the MC is working again as before the shutdown (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

The (new) AA of the IB is ready to be implemented: it should replace the LC of the IB

IMC automatic alignment


Power stabilization
Power Stabilization (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

IMC_trans HF

Master laser

Diode current

Data from spectrum analyser is 50 times lower in the

100Hz region than dataDisplay measurent (above) (to be investigated)


Rfc automatic alignment

Closed on the vertical DOF (1 Hz bandwidth) (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

Horizontal DOF not working (M16_PZT horiz. not working properly)

RFC automatic alignment

M15

RFC AA now

separated from IMC AA

M16

Wavefront sensing

quadrants on the EIB


New pr mirror monolithic and higher r 0 95 vs 0 925
New PR mirror: monolithic and higher R (0.95 vs 0.925) (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

Comparison between the old (black) and new monolithic (red) PR mirror


Virgo commissioning progress

II- ITF restart (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.


Ib lc after rfc locking
IB LC after RFC locking (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

The IB LC filters were changed after restart: frequency noise was reintroduced


Frequency noise after rfc locking itf misaligned
Frequency noise after RFC locking (ITF misaligned) (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

IB transfer function

Frequency noise


Ib lc filter upgrade
IB LC filter upgrade (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

IB in high noise control

IB in low noise control

Blue curves: MC in high/low noise (factor 100 G difference)

Red curves: MC in high/low noise (factor 100 G difference)


Mc rfc locking filter after restart
MC-RFC locking filter: after restart (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

Too small G  SSFS couldn’t engage


Frequency noise after rfc locking itf misaligned1
Frequency noise after RFC locking (ITF misaligned) (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

Purple: C7

Black: after new IB RFC locking


New mc rfc locking filter upgrade
New MC-RFC locking filter upgrade (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.


New frequency noise ssfs can engage
New frequency noise: SSFS can engage (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.


Frequency noise c7 beam attenuated vs now
Frequency noise: C7 (beam attenuated) vs now (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom.

Now

C7


Itf restart locking

Photodiodes gains readjusted for the higher power (10 times more than before shutdown)

Automatic alignment restarted:

Tested the drift control of the two cavities

Quadrant photodiodes gains adjusted

Not gone further because it needs the full locking of the ITF

PR-NI

PR-NE

CITF

NE-WE-BS

NA

WA

Recombined

SSFS

Variable Finesse with PR misaligned

….

Now the full recycled ITF is being locked

ITF restart: locking

Since end of January:


Virgo commissioning progress
SSFS more than before shutdown)

L1

L2

B2_3f_ACp

B5_ACp

SSFS using CARM for the

Beam Frequency stabilization

B1p_DC

B8p

B2_3f_ACp  PR (PRCL)

B1p_DC  BS (MICH)

B5_ACp  SSFS (CARM)

B8_ACp NE-WE (DARM)

l1

l2

l0


Ssfs restart

The restart of SSFS required some gain tuning because: more than before shutdown)

RFC locking filter was different

Frequency noise was different

RFC gains are different

SSFS restart

After tuning readjustment, IB LC filters and RFC locking filter SSFS is now working properly


Towards the dark fringe
Towards the dark fringe more than before shutdown)

B8p

L1

L2

B2_3f_ACp

B5_ACp

SSFS using CARM

STEP 7

B2_3f_ACp  PR (PRCL)

B1p_DC  BS (MICH)

B5_ACp  SSFS (CARM)

B8_ACp NE-WE (DARM)

l1

l2

l0

B1p_DC

DF 0.05



Parabolic telescope t2 m6 mirror misalignment
Parabolic telescope T2: M6 mirror misalignment dark fringe)

3 closed loop picomotors

Tilted: tx (650 mrad) and ty (850 mrad)


Effect of m6 misalignment
Effect of M6 misalignment dark fringe)

M6 aligned (beam at 3 km)

M6 misaligned (beam at 3 km)

(Qx=650 mrad

Qy=850 mrad)


Resulting beam shape

B dark fringe)eam size at NE:

wx= 60mm

wy= 60mm

Beam size on Cam5:

wx= 0.7mm

wy= 0.42mm

To fit these measurements, the beam parameters inside ITF with respect to BS position are(perfect matching for w=21.3, z=0):

wx= 15mm, zx= -400m

wy= 23mm, zy= 800m

The matching on the FP cavities should be around89% in theseconditions: confirmed by measurement (88%).

Beam parameters before shutdown:

wx= 18mm, zx= -500m

wy= 18mm, zy= 500m)

Resulting beam shape


Short telescope t1 centering
Short telescope T1 centering dark fringe)

DY

L1

f=130

L2

f=-80

D=60mm

ccd


Beam shape after the faraday isolator
Beam shape after the Faraday isolator dark fringe)

Dy=4 mm Dw/w = 9%

Dy=2 mm Dw/w = 2%


Conclusion

The restart of the Injection System has required more than two months

The new Injection System is now working

The ITF is close to final locking with full power (8 W)

Some problems involving the new IB remain open:

The RFC matching is poor and the RFC transmission is not sent out of the IB tower

The astigmatism of the beam has to be corrected

Conclusion

For the moment the effort is in achieve the full locking again and, since we can go on in these conditions, to pospone the solving of this problems to a later moment


ad
  • Login